# Periodic Asynchrony: An On-Policy Approach for Accelerating LLM Reinforcement Learning

### v4

@@ 6 Experiments
6.1 Model and Environment Configuration
The models compared in this section are Qwen2.5-7B-Instruct Yang et al. (2024), Qwen3-8B Team (2025), and DeepSeek-R1-Distill-Qwen-32B DeepSeek-AI (2025), as the current version of MindSpeed-RL does not yet support reinforcement learning training for Qwen3-32B. The training data comes from the publicly available math problem datasets DeepScaler Luo and others (2025) and GSM8K Cobbe et al. (2021). All reinforcement learning experiments use GRPO algorithm. The accuracy test set is the AIME24 Zhang and Math-AI (2024).
6.2 Asynchronous Training Speed Comparison
Since the prompts in DeepScaler are significantly shorter than the responses, Shared-Prompt Attention is disabled; nevertheless, our method still achieves substantial performance improvements. In the first set of experiments, Qwen3-8B was used as the training model, with a context length of 16k. As shown in Table 1, in terms of TPSPD, the training speed ratio among the four implementations was $\mathbf{1\!:\!2.52\!:\!1.62\!:\!3.12}$. The model used in the second set of experiments is DeepSeek-R1-Distill-Qwen-32B. As shown in Table 2, the evaluation methodology is identical to that used in the first set of experiments. According to the experimental results, for the 32B-scale model, the TPSPD ratio between MindSpeed-RL and our asynchronous framework reaches $\mathbf{1\!:\!5.05}$ under a 16K context length. Due to consistent out-of-memory issues with VERL at 16K, we additionally evaluate VERL under an 8K context length, where the TPSPD ratio between VERL and our asynchronous framework is $\mathbf{1\!:\!1.76}$. This experiment also shows that insufficient context-length configuration can lead to a degradation in model accuracy. We further observe that, under this regime, inference length is a primary factor affecting end-to-end throughput.
Given the increased problem difficulty, DeepScaler represents a regime in which inference is more time-consuming than training. To study the complementary setting where training dominates inference, we additionally conduct experiments on GSM8K using Qwen2.7-7B-Instruct, where both prompts and responses are relatively short. We compare MindSpeed-RL, VERL, and our asynchronous framework under configurations with Shared-Prompt Attention enabled and disabled, and evaluate their end-to-end throughput and accuracy. As shown in Table 3, Shared-Prompt Attention significantly reduces the number of training tokens and leads to improved training throughput, with the corresponding TPSPD ratios of $\mathbf{1\!:0.91\!:\!0.27\!:\!2.19}$.
@@ 6.2 Asynchronous Training Speed Comparison
Async (ours) 0.737 48 32 1 123.999 M 33.449
@@
Table 2: All frameworks are trained for 20 steps. For GRPO, each group performs 32 rollouts. In the additional group of experiments, the global batch size (GBS) is increased to 64, since VERL requires a minimum micro-batch size (MBS) of 64 when running on 64 NPUs.
@@
Async (ours) 0.912 1 84.192 M 54.463
Async (ours) 0.919 16 60.926 M 435.596
@@
Table 3: All frameworks are trained on 16 NPUs and complete training on the full GSM8K training set at step 116. For GRPO, each group performs 32 rollouts, and the context length is fixed to 1K for all settings. In our asynchronous framework, a micro-batch size of 1 indicates that Shared-Prompt Attention is disabled, whereas a micro-batch size of 16 means that 16 rollouts share a single prompt. Evaluation is conducted on the GSM8K test set.
For accuracy evaluation, we employed a rule-based method where the predicted answer from the policy model is considered correct if it can be accurately extracted and matches the ground-truth answer in the dataset; otherwise, it is deemed incorrect. The tests were conducted using the inference en gine vLLM 0.9.1.rc. The results show that the models trained by all four frameworks achieved similar accuracy, with any differences falling within the expected range of random fluctuations. This supports the equivalence in performance between the asynchronous and synchronous methods implemented in this study.
We also monitored the reward scores during training for all four frameworks, as shown in Figure 5. The reward trajectories of our synchronous and asynchronous methods almost completely overlap, suggesting that the asynchronous approach is computationally equivalent to the synchronous one. Both curves exhibit a slight upward trend, likely because the baseline model was pre-trained on similar data.
Figure 5: Training step-wise average reward score comparison.
6.3 Asynchronous Scalability Experiments
@@
We also conducted a set of experiments to evaluate the scalability of our framework, using the same configuration as in the first experiment group. The results are shown in Table 4.
@@
When expressed in TPS (tokens per second), the results are more intuitively illustrated in Figure 6. Training with 32 NPUs shows a 1.83× speedup compared to 16 NPUs, while using 64 NPUs results in a 1.9× speedup over 32 NPUs. Overall, the proposed asynchronous reinforcement learning framework demonstrates near-linear scaling in training speed as the number of devices increases.
@@ 7 Conclusion
This paper introduces an asynchronous on-policy reinforcement learning algorithm that maintains the same training stability as synchronous RL methods. The feasibility of performing reinforcement learning training with micro-batches is theoretically demonstrated. Prompt-level asynchronous parallelism is employed to maximize the overlap between inference and training. A unified three-mode (policy, old policy, reference) architecture with identically distributed components is introduced, along with a shared-prompt attention mechanism that offers distinct advantages for long prompts and short responses. Experimental results demonstrate that the proposed framework achieves a three- to five-fold improvement in end-to-end performance over current mainstream RL training frameworks on the NPU platform. These results further support the equivalence in training effectiveness between the asynchronous and synchronous methods. Additionally, the decoupling of training and inference enables independent and flexible scaling, optimizing the balance between speed and device utilization. The decoupling of the training architecture and algorithm also ensures the proposed framework’s compatibility with other on-policy reinforcement learning algorithms, allowing them to benefit from the advantages of asynchronous training without requiring algorithmic modifications, nor ensuring the stability of modified algorithms.

### v5

@@ 6.1 Model and Environment Configuration
Figure 5: Training step-wise average reward score comparison.
6.2 Training Throughput Comparison
6.2.1 Comparison with Existing Frameworks
8B Model on DeepScaleR.
Table 1 reports results for Qwen3-8B trained on DeepScaleR with a 16K context length. Since prompts in this dataset are significantly shorter than responses, Shared-Prompt Attention is disabled for all methods in this experiment. Our asynchronous framework achieves a TPSPD of 192.259, outperforming MindSpeed-RL ($\mathbf{3.12\times}$), VERL ($\mathbf{1.24\times}$), and our synchronous baseline ($\mathbf{1.92\times}$). The $1.92\times$ speedup over the synchronous baseline is particularly noteworthy, as it closely approaches the theoretical upper bound of $2\times$ derived in Section 4.2, validating the efficiency analysis of the periodic asynchronous framework in a practical setting.
32B Model on DeepScaleR.
Table 2 reports results for DeepSeek-R1-Distill-Qwen-32B. In the first group, our asynchronous framework running on only 48 NPUs achieves a TPSPD of 33.449, delivering a $\mathbf{5.05\times}$ speedup over MindSpeed-RL running on 64 NPUs. This result is particularly significant: not only does our method achieve substantially higher throughput, it does so with fewer hardware resources, demonstrating both efficiency and resource economy. In the second group, all frameworks are evaluated at 8K context to align with VERL, which encounters out-of-memory issues at 16K. VERL’s accuracy drops sharply to 0.617—well below the base model level—while our framework maintains 0.675, a degradation we attribute to response truncation at insufficient context length. Our asynchronous framework achieves a further $\mathbf{1.76\times}$ speedup over VERL even under these favorable conditions for VERL.
7B Model on GSM8K.
Table 3 studies the training-dominated regime using Qwen2.5-7B-Instruct on GSM8K with a 1K context. Our asynchronous framework with Shared-Prompt Attention enabled achieves a TPSPD of 435.596, the highest among all settings, corresponding to a $\mathbf{2.19\times}$ speedup over MindSpeed-RL and $\mathbf{2.39\times}$ over VERL. Notably, both competing frameworks also achieve competitive accuracy, confirming that the throughput gap is not due to shortcuts in training quality.
@@ 6.2.1 Comparison with Existing Frameworks
Sync (ours) 0.725 48 32 1 123.613 M 26.219
Async (ours) 0.738 48 32 1 123.999 M 33.449
@@
Table 2: All frameworks are trained for 20 steps. For GRPO, each group performs 32 rollouts. In the additional group of experiments, the global batch size (GBS) is increased to 64, since VERL requires a minimum micro-batch size (MBS) of 64 when running on 64 NPUs. VERL is evaluated at 8K context due to out-of-memory issues at 16K.
@@
Async (ours), w/o SPA 0.912 1 84.192 M 54.463
Sync (ours), w/ SPA 0.930 16 60.586 M 218.396
Async (ours), w/ SPA 0.919 16 60.926 M 435.596
@@
Table 3: All frameworks are trained on 16 NPUs and complete training on the full GSM8K training set at step 116. For GRPO, each group performs 32 rollouts, and the context length is fixed to 1K for all settings. SPA denotes Shared-Prompt Attention. A micro-batch size of 1 indicates SPA is disabled, whereas 16 means 16 rollouts share a single prompt. Evaluation is conducted on the GSM8K test set.
6.2.2 Architectural Analysis
To isolate the contribution of our architectural design from the benefit of asynchronous execution, we compare Sync (ours) directly against MindSpeed-RL, the strongest competing synchronous framework. Under identical synchronous execution, our framework achieves a $\mathbf{1.62\times}$ speedup on the 8B model (Table 1) and a $\mathbf{3.96\times}$ speedup on the 32B model (Table 2), attributable to the efficiency of the decoupled training–inference design and the unified tri-model architecture. Furthermore, in the training-dominated GSM8K setting, Sync (ours), w/ SPA (TPSPD: 218.396) already surpasses VERL (async) (TPSPD: 182.560) despite using synchronous execution, demonstrating that Shared-Prompt Attention alone is sufficient to overcome the throughput gap in training-dominated regimes. These results confirm that the architectural contributions of this work provide independent and meaningful efficiency gains beyond asynchronous execution alone.
6.2.3 Ablation Analysis
Table 3 enables a clean ablation of the two key components proposed in this work: the periodic asynchronous framework and Shared-Prompt Attention. We isolate their respective contributions as follows.
Effect of Shared-Prompt Attention.
Comparing Async (ours), w/o SPA (TPSPD: 54.463) with Async (ours), w/ SPA (TPSPD: 435.596) under otherwise identical settings reveals that enabling Shared-Prompt Attention alone yields an $\mathbf{8\times}$ improvement in throughput. This gain stems from two sources: a reduction in training tokens from 84.192 M to 60.926 M due to shared prompt computation, and the elimination of padding overhead within micro-batches. These results are consistent with the $K$-fold complexity reduction predicted by the analysis in Section 4.3, where $K=16$ rollouts share a single prompt.
Effect of Periodic Asynchrony.
Comparing Sync (ours), w/ SPA (TPSPD: 218.396) with Async (ours), w/ SPA (TPSPD: 435.596) isolates the contribution of asynchronous execution under identical architecture and data conditions. The asynchronous framework delivers a $\mathbf{2\times}$ speedup, closely matching the theoretical upper bound established in Section 4.2, and confirming that the overlap between inference and training is effectively exploited in practice.
Combined Effect.
The two components are complementary in nature: Shared-Prompt Attention reduces the per-step training cost, while periodic asynchrony minimizes idle waiting time between inference and training. Together, they deliver a $\mathbf{2.19\times}$ speedup over the strongest competing framework MindSpeed-RL and an $\mathbf{8\times}$ speedup over our unoptimized baseline (Async w/o SPA), demonstrating that the benefits of the two components are largely multiplicative.
6.2.4 Accuracy Validation
For accuracy evaluation, we employ a rule-based method where the predicted answer is considered correct if it can be accurately extracted and matches the ground-truth answer in the dataset; otherwise it is deemed incorrect. As shown across all tables, the accuracy of our framework remains consistent with competing methods, with differences falling within the expected range of random fluctuation. Furthermore, the reward trajectories of our synchronous and asynchronous methods almost completely overlap throughout training, as shown in Figure 5, empirically corroborating the theoretical equivalence established in Theorem 1. These results confirm that the substantial throughput gains introduced by our framework come at no cost to training effectiveness.
6.3 Scalability Analysis
@@
We conduct a set of experiments to evaluate the scalability of our framework using the same configuration as the first experiment group. The results are shown in Table 4.
@@
When expressed in TPS (tokens per second), the results are more intuitively illustrated in Figure 6. Training with 32 NPUs shows a $1.83\times$ speedup compared to 16 NPUs, while using 64 NPUs results in a $1.9\times$ speedup over 32 NPUs, demonstrating near-linear scaling in total throughput. The moderate decrease in TPSPD with increasing device count is expected, as inter-node communication overhead grows with scale while the training-to-inference ratio is held fixed at 1:4. Overall, the proposed asynchronous reinforcement learning framework demonstrates strong scalability as the number of devices increases.
@@ 7 Conclusion
This paper addresses the training efficiency bottleneck in on-policy reinforcement learning by proposing a periodically asynchronous framework that achieves provably equivalent updates to synchronous training. By introducing a temporary data generator between the data loader and the trainer, the framework transforms synchronous RL into an asynchronous producer–consumer pipeline, maximizing the overlap between inference and training without any algorithmic modifications. We theoretically establish that micro-batch training is fully compatible with on-policy RL objectives, and prove that neither the on-policy correctness nor the gradient equivalence is compromised by asynchronous execution. A unified tri-model architecture with identically distributed components is introduced to support efficient and consistent weight synchronization, along with a shared-prompt attention mechanism that significantly reduces redundant computation in long-prompt, short-response settings. Experimental results on NPU platforms demonstrate a three- to five-fold improvement in end-to-end throughput over mainstream RL frameworks, while maintaining fully comparable training effectiveness. Furthermore, the decoupling of training and inference enables independent and flexible scaling, and the algorithm-agnostic design ensures compatibility with any on-policy RL method, allowing the community to benefit from asynchronous acceleration without sacrificing theoretical guarantees.